Maria is 16 years younger than Ashley. If the sum of their ages is 30, how old is Maria ?
Answer by stanbon(75887)

Maria is 16 years younger than Ashley. If the sum of their ages is 30, how old is Maria ?
------
Equations:
m = a-16
m + a = 30
---------------------
Rearrange:
m-a = -16
m+a = 30
-----
Add to get::
2m = 14
Ans: Maria is 7 yrs old
-------------
Cheers,
Stan H.
------------